How many moles of Pb are in 4.71 × 1021 Pb atoms?
1
Expert's answer
2018-03-27T08:00:16-0400
Solution:
1 mole = 6.022 ∙ 10^23 atoms, so:
Amount of Pb: (4.71 ∙ 10^21) / (6.022 ∙ 10^23) = 0.782 ∙ 10-2 = 7.82 ∙ 10-3 = 0.00782 mol
Answer:
7.82 ∙ 10-3 mol = 0.00782 mol